Question 17 5 pts The following will be used to answer the next question Debt: 15,000 10% coupon bonds outstanding, 30 yearsQuestion 18 5 pts The following will be used to answer the next question Debt: 15,000 10% coupon bonds outstanding. 30 yearsQuestion 19 5 pts The following will be used to answer the next question Debt: 15,000 10% coupon bonds outstanding, 30 years
Answer #1

Q17

cost of debt can be found from excel using " =2*rate(60,50,-1060,1000)" formula ;

cost of debt =9.4%

Q18:

cost of equity = risk free rate +beta*market risk premium =6+1.5*7.5 =17.25%

Q19:

cost of preferred stock =7/128 =5.5%

Q20:

market value of Debt=1060*15000=15900000

market value Equity =300000*80=24000000

market value preferred share=20000*128=2560000

weight of equity =24000000/(15900000+24000000+2560000) =57%